# MLLM-FL: Multimodal Large Language Model Assisted Federated Learning on Heterogeneous and Long-tailed Data

### v1

@@ Abstract.
Previous studies on federated learning (FL) often encounter performance degradation due to data heterogeneity among different clients. In light of the recent advances in multimodal large language models (MLLMs), such as GPT-4v and LLaVA, which demonstrate their exceptional proficiency in multimodal tasks, such as image captioning and multimodal question answering. We introduce a novel federated learning framework, named Multimodal Large Language Model Assisted Federated Learning (MLLM-FL), which which employs powerful MLLMs at the server end to address the heterogeneous and long-tailed challenges. Owing to the advanced cross-modality representation capabilities and the extensive open-vocabulary prior knowledge of MLLMs, our framework is adept at harnessing the extensive, yet previously underexploited, open-source data accessible from websites and powerful server-side computational resources. Hence, the MLLM-FL not only enhances the performance but also avoids increasing the risk of privacy leakage and the computational burden on local devices, distinguishing it from prior methodologies. Our framework has three key stages. Initially, prior to local training on local datasets of clients, we conduct global visual-text pretraining of the model. This pretraining is facilitated by utilizing the extensive open-source data available online, with the assistance of multimodal large language models. Subsequently, the pretrained model is distributed among various clients for local training. Finally, once the locally trained models are transmitted back to the server, a global alignment is carried out under the supervision of MLLMs to further enhance the performance. Experimental evaluations on established benchmarks, show that our framework delivers promising performance in the typical scenarios with data heterogeneity and long-tail distribution across different clients in FL.
Federated Learning, Multimodality, Large Language Model
copyright: nonejournalyear: conference: ; ;
1. Introduction
The surge in IoT devices has unlocked vast potential for leveraging edge-generated data in driving cooperative computing applications such as autonomous vehicles, video analytics, and recommendation systems. Traditionally, the centralized training process raises significant data privacy and security concerns due to the necessity of transferring local information. Federated learning (FL), as introduced by (McMahan et al., 2017), offers a solution to these privacy challenges by enabling collaborative model training across numerous clients under the orchestration of a central server, without sharing the raw data. FL systems bring obvious advantages by involving clients downloading a global model, performing local updates using their data, and then sending these updates back to the server. The server aggregates these updates to enhance the global model, thereby preserving data privacy. Aside from the privacy considerations mentioned above, there are two fundamental acknowledgements about (cross-device) federated learning which have been widely recognized: data heterogeneity among clients and the limited and diverse computational resources on local devices (McMahan et al., 2017; Kairouz et al., 2021; Mahlool and Abed, 2022; Zhang et al., 2024).
Data heterogeneity represents a significant challenge in federated learning. It largely stems from the fact that the data across participating clients are distributed independently, with each client having a different sample distribution. Due to the diversity in clients’ datasets. these datasets often exhibit a long-tailed distribution, leading to client models that are biased toward the more common classes (Sahu et al., 2018; Wang et al., 2020b). This discrepancy often results in a drop in model accuracy. Although several approaches have been proposed (Li et al., 2022; Huang et al., 2021, 2023; Feng et al., 2023; Shi et al., 2023; Jia et al., 2023; Yang et al., 2021; Hao et al., 2021), the majority fail to strike an optimal balance between performance and mitigating two critical issues: 1. avoiding concerns of privacy leakage, and 2. preventing the imposition of extra computational loads on local edge devices. For instance, some contemporary methodologies necessitate the transmission of both gradients and parameters from local models to the server, which introduces substantial privacy risks. This is because attackers could potentially reverse-engineer the transmitted data to reconstruct client-specific images, as highlighted in various studies (Geiping et al., 2020; Zhu et al., 2019; Haim et al., 2022). Alternatively, other approaches require the deployment of sizable models on local devices, which increases the memory and computational demands.
@@ 1. Introduction
Figure 1. The whole workflow of MLLM-FL The MLLM are utilized in the first stage Global Multimodal Pretraining and the third stage Global Alignment on the server side, to avoid extra computational load on devices.
@@
In light of the current popularity and exceptional proficiency of multimodal large language models (MLLMs) in tasks involving multimodalities, such as image captioning and multimodal question answering (OpenAI, 2022; Team et al., 2023; Li et al., 2023b; Liu et al., 2023; Zhu et al., 2023), we introduce a three-stage framework, named Multimodal Large Language Model sisted Federated Learning (MLLM-FL), which utilizes multimodal large language models (MLLMs) to the FL performance on heterogeneous and long-tailed data. The adaptation of MLLMs in FL is supported by two main considerations. Firstly, beyond the heterogenous and long-tailed distribution of client datasets, there exists an abundance of open-sourced and legally available data on the internet that can be utilized for training. This implies the potential of employing MLLMs to annotate unstructured, unlabeled online data, thereby augmenting FL performance. Secondly, in contrast to the limited computational resources available on client devices, the server-side capabilities are significantly more robust. This disparity opens up the possibility of deploying additional, more powerful MLLM on the server side to provide assistance to the FL system.
Table 1. Comparison between our methods and other status quo approaches for addressing long-tailed distribution challenges in FL
@@
CReFF (Shang et al., 2022),
CLIP2FL (Shi et al., 2023)
MLLM-FL (ours)
@@
  • Firstly, we pioneer the integration of the widely recognized multimodal large language model (MLLM) as an auxiliary tool in federated learning, aiming to enhance the utilization of previously underexplored internet data resources and server computational capabilities. Leveraging the formidable cross-modality representation capabilities and the vast open-vocabulary prior knowledge inherent in MLLMs, we introduce novel methodologies to address the challenges posed by long-tail distributions and data heterogeneity. This also marks the first exploration of employing MLLMs to augment federated learning, establishing an innovative framework in the field.
@@
  • Our extensive experimental results show that MLLM-FL can effectively handle heterogeneity and class-distribution imbalance, consistently surpassing the performance of existing state-of-the-art federated learning methodologies across a variety of datasets.
2. Related Work
2.1. Multimodal large language model
@@
The introduction of GPT-4(Vision) (OpenAI, 2022) and Gemini (Team et al., 2023) have demonstrated remarkable abilities in Multimodal understanding and generation, sparking a research fervor on Multimodal large language model. This enthusiasm extends to a variety of tasks, including image-text comprehension (Li et al., 2023b; Liu et al., 2023; Zhu et al., 2023); video-text understanding (Li et al., 2023a; Maaz et al., 2023); and audio-text understanding (Chu et al., 2023). Among them, recent studies in image-text comprehension with large vision-language models (LVLM) (Li et al., 2023b; Liu et al., 2023; Zhu et al., 2023) have catalyzed notable advancements in harnessing the robust capabilities of large language models to tackle multimodal tasks effectively, such as crafting narratives from images and executing intricate reasoning tasks. Prominent instances include Visual ChatGPT (Wu et al., 2023), which amalgamates diverse visual foundational models for intricate visual tasks and instructions, employing iterative feedback to synchronize visual and textual modalities. In a similar way, MM-REACT (Yang et al., 2023) merges ChatGPT with visual models for multimodal undertakings, especially in the Visual Question Answering (VQA) framework. BLIP-2 (Li et al., 2023b), notable for its Q-former model, has shown encouraging outcomes in VQA tasks, both in zero-shot and fine-tuning settings. LLaMA-Adapter (Gao et al., 2023) enhances multimodal fine-tuning efficiency by integrating adaptation prompt vectors as adjustable parameters, showcasing versatility in multimodal contexts. MiniGPT-4 (Zhu et al., 2023), derived from GPT-4 and incorporating elements from BLIP-2 and Vicuna (Chiang et al., 2023), specializes in caption generation and model refinement through image-text pair fine-tuning. LLaVA (Liu et al., 2023), leveraging GPT-4, focuses on a broad spectrum of instruction fine-tuning data, ranging from multi-turn QA to image descriptions, adopting a dual-stage fine-tuning approach that prioritizes language model loss while keeping the visual model static.
2.2. Federated Learning with Heterogeneous Data
@@
Current methodologies tackling the challenge of data heterogeneity fall into the following broad categories. Some approaches aim to simultaneously improve the models on both clients and server sides through optimization techniques. Key contributions in this area have been made by the work of (Li et al., 2022; Huang et al., 2021, 2023; Feng et al., 2023), who have investigated various optimization methods. Other strategies focus on enhancing the stability of local models via knowledge transfer, a technique that is model-agnostic and has been explored in the research by (Xu et al., 2023; Zhang et al., 2021; Chen et al., 2020), aiming to mitigate data heterogeneity by spreading local training knowledge throughout the whole FL framework. Additional methods, such as those proposed by (Lin et al., 2020; Chen and Chao, 2020), concentrate on improving model aggregation on the server side to address data heterogeneity. Certain strategies also regulate the scheduling of client participation to avoid biasing the FL model towards classes that are more prevalent, as explored in the studies by (Yang et al., 2020; Zhang et al., 2023a). While these approaches have advanced the handling of data heterogeneity, they often do not fully address the specific issues related to long-tailed distributions in FL.
Figure 2. The visulalization of our pretraining mechanism
@@
A recent approach, CReFF (Shang et al., 2022), introduces a decoupling strategy to create balanced class-distribution federated features for the server model and to retrain the classifier with these features. Nonetheless, CReFF encounters two main limitations due to its reliance on generating federated features through client-side gradient information: 1) The one-to-many relationship between gradients and samples can result in the problem becoming ill-posed; 2) The absence of semantic guidance might lead to federated features that lack discriminative ability for their respective classes. The subsequent attempt, CLIP2FL (Shi et al., 2023), seeks to overcome these drawbacks by integrating a multimodal model to direct the federated learning process. However, it still has its own drawbacks. Firstly, deploying the sizable CLIP model on devices increases memory and computational demands. Secondly, transmitting both the gradient and parameters of local models to the server, as necessitated by both CLIP4FL and CReFF, raises significant privacy concerns, as attackers could potentially reconstruct client images through reverse engineering (Geiping et al., 2020; Zhu et al., 2019; Haim et al., 2022).
3. Methodology
@@
In the conventional federated learning (FL) pipeline, the FL models are typically assigned to local clients for training on their heterogeneous datasets. These models are then sent back to the server for aggregation. This cycle continues repeatedly until the FL training concludes. To better align the above FL framework with practical requirements, we incorporate the following two additional stages. The first stage occurs before local training, involving pretraining on the server side, a strategy supported by previous work (Chen et al., 2022) which found that pretraining can accelerate the convergence of FL training and mitigate the effects of data heterogeneity on convergence. The second stage takes place after aggregation, where the FL model may undergo further training to meet the broader requirements of FL companies, such as the performance and safety considerations we discuss later.
@@
Drawing inspiration from recent developments in learning paradigms of natural language processing, we structure our work into three parts: global multimodal pretraining, federated local finetuning, and global alignment. We deploy multimodal approaches at the server side to assist both global multimodal pretraining and global alignment phases. In this section, we will introduce our comprehensive framework as follows: Section 3.1 will delve into the details of our global multimodal pretraining; Section 3.2 will cover federated local finetuning; Section 3.3 will discuss global alignment; and in Section 3.4, we will compare our method with previous approaches.
3.1. Global Multimodal Pretraining
@@
As discussed in Section 1, the wealth of open-source multimodal data, such as images and their captions, remains underutilized resources for pretraining FL models. Often, these datasets are noisy, unlabeled or contain elements that are too complex, making them unsuitable for straightforward pretraining of the compact FL models. However, given the current advanced capabilities in multimodal processing of MLLMs, we now have new, convenient methods to leverage such data for pretraining purposes. Utilizing GPT-4, akin to the approach used in LLaVA, we can transform complex image data collected from the internet into three main categories:
@@
Our pretraining mechanism draws inspiration from the structure of LLaVA, a highly effective and recent multimodal large language model. It consists of three key components: a visual encoder $g$, which is a frozen pretrained CLIP model; a projection layer designed to align the features of the visual model with the text domain embeddings, where the projection layer is a trainable matrix $W$; and a part comprising a large language model (LLM), typically employing promising models such as Vicuna or LLaMA-2. The workflow of LLaVA proceeds as follows: For the data formats mentioned earlier, whether it be conversation, detailed description, or complex reasoning, the input includes a text modality instruction $\mathbf{X}_{\mathrm{q}}$ (e.g., ”Could you provide a detailed description of this image?”) and an image $\mathbf{X}_{\mathrm{v}}$. The instruction $\mathbf{X}_{\mathrm{q}}$ passes through an embedding layer to obtain the text embedding $\mathbf{H}_{q}$. As for the image $\mathbf{X}_{\mathrm{v}}$, it first goes through the visual encoder $g$ to acquire the grid feature $\mathbf{Z}_{\mathrm{v}}$, which then passes through the projection layer to obtain the visual embedding $\mathbf{H}_{v}$, aligned in dimension with $\mathbf{H}_{q}$:
$\mathbf{H}_{v}=\mathbf{W}\cdot\mathbf{Z}_{\mathrm{v}}\text{, where }\mathbf{Z}_{\mathrm{v}}=g\left(\mathbf{X}_{\mathrm{v}}\right)$
@@
Drawing from the LLaVA mechanism, our Global Multimodal Pretraining essentially integrates the compact FL model, designed for downstream image classification tasks, as a component within the LLaVA framework’s visual encoder. The FL model is made trainable and is denoted as $g_{fl}$. Inspried by the previous work in knowledge distillation (Hinton et al., 2015; Mirzadeh et al., 2020; Jin et al., 2019; Park et al., 2019; Zhang et al., 2023b), We have developed an approach termed Dynamic Weighted Distillation, which involves computing a weighted average of the visual features obtained from the FL model and those from the original visual encoder:
$\mathbf{Z}_{\mathrm{v}}=(1-\alpha)g\left(\mathbf{X}_{\mathrm{v}}\right)+\alpha g_{fl}\left(\mathbf{X}_{\mathrm{v}}\right)$

### v2

# MLLM-LLaVA-FL: Multimodal Large Language Model Assisted Federated Learning

@@ Abstract
Previous studies on federated learning (FL) often encounter performance degradation due to data heterogeneity among different clients. In light of the recent advances in multimodal large language models (MLLMs), such as GPT-4v and LLaVA, which demonstrate their exceptional proficiency in multimodal tasks, such as image captioning and multimodal question answering. We introduce a novel federated learning framework, named Multimodal Large Language Model Assisted Federated Learning (MLLM-LLaVA-FL), which employs powerful MLLMs at the server end to address the heterogeneous and long-tailed challenges. Owing to the advanced cross-modality representation capabilities and the extensive open-vocabulary prior knowledge of MLLMs, our framework is adept at harnessing the extensive, yet previously underexploited, open-source data accessible from websites and powerful server-side computational resources. Hence, the MLLM-LLaVA-FL not only enhances the performance but also avoids increasing the risk of privacy leakage and the computational burden on local devices, distinguishing it from prior methodologies. Our framework has three key stages. Initially, we conduct global visual-text pretraining of the model. This pretraining is facilitated by utilizing the extensive open-source data available online, with the assistance of MLLMs. Subsequently, the pretrained model is distributed among various clients for local training. Finally, once the locally trained models are transmitted back to the server, a global alignment is carried out under the supervision of MLLMs to further enhance the performance. Experimental evaluations on established benchmarks, show that our framework delivers promising performance in the typical scenarios with data heterogeneity and long-tail distribution across different clients in FL.
1 Introduction
@@
Figure 1: The workflow of MLLM-LLaVA-FL. The MLLM are utilized in the first stage Global Multimodal Pretraining and the third stage Global Alignment on the server side, to avoid extra computational load on devices.
Data heterogeneity represents a significant challenge in federated learning[37, 22, 42, 66, 14, 60]. It largely stems from the fact that the data across participating clients are distributed independently, with each client having a different sample distribution. Due to the diversity in clients’ datasets, these datasets often exhibit a long-tailed distribution, leading to client models that are biased toward the more common classes [42, 51]. This discrepancy often results in a drop in model accuracy. Although several approaches have been proposed [26, 17, 18, 10, 45, 57, 19, 9], the majority fail to strike an optimal balance between performance and mitigating two critical issues: 1. avoiding concerns of privacy leakage, and 2. preventing the imposition of extra computational loads on local edge devices. For instance, some contemporary methodologies necessitate the transmission of both gradients and parameters from local models to the server, which introduces substantial privacy risks. This is because attackers could potentially reverse-engineer the transmitted data to reconstruct client-specific images, as highlighted in various studies [12, 72, 13]. Alternatively, other approaches require the deployment of sizable models on local devices, which increases the memory and computational demands.
@@
In light of the current popularity and exceptional proficiency of multimodal large language models (MLLMs) in tasks involving multimodalities, such as image captioning and multimodal question answering, we introduce a three-stage framework, named Multimodal Large Language Model sisted Federated Learning (MLLM-LLaVA-FL), which utilizes multimodal large language models (MLLMs) to the FL performance on heterogeneous and long-tailed data. The adaptation of MLLMs in FL is supported by two main considerations. Firstly, beyond the heterogenous and long-tailed distribution of client datasets, there exists an abundance of open-sourced and legally available data on the internet that can be utilized for training. This implies the potential of employing MLLMs to annotate unstructured, unlabeled online data, thereby augmenting FL performance. Secondly, in contrast to the limited computational resources available on client devices, the server-side capabilities are significantly more robust. This disparity opens up the possibility of deploying additional, more powerful MLLM on the server side to provide assistance to the FL system.
@@
CReFF [44],
CLIP2FL [45]
MLLM-LLaVA-FL (ours)
Table 1: Comparison between our methods and other status quo approaches for addressing long-tailed distribution challenges in FL
@@
  • Firstly, we pioneer the integration of the widely recognized multimodal large language model (MLLM) as an auxiliary tool in federated learning, aiming to enhance the utilization of previously underexplored internet data resources and server computational capabilities. Leveraging the formidable cross-modality representation capabilities and the vast open-vocabulary prior knowledge inherent in MLLMs, we introduce novel methodologies to address the challenges posed by long-tail distributions and data heterogeneity.
@@
  • Our extensive experimental results show that MLLM-LLaVA-FL can effectively handle heterogeneity and class-distribution imbalance, consistently surpassing the performance of existing state-of-the-art federated learning methodologies across a variety of datasets.
2 Related Work
2.1 Multimodal large language model
@@
The introduction of GPT-4(Vision) [39] and Gemini [48] have demonstrated remarkable abilities in Multimodal understanding and generation, sparking a research fervor on Multimodal large language model. This enthusiasm extends to a variety of tasks, including image-text comprehension [27, 31, 71]; video-text understanding [28, 35]; and audio-text understanding [6]. Among them, recent studies in image-text comprehension with large vision-language models (LVLM) [27, 31, 71] have catalyzed notable advancements in harnessing the robust capabilities of large language models to tackle multimodal tasks effectively, such as crafting narratives from images and executing intricate reasoning tasks. Prominent instances include Visual ChatGPT [54], which amalgamates diverse visual foundational models for intricate visual tasks and instructions, employing iterative feedback to synchronize visual and textual modalities. In a similar way, MM-REACT [58] merges ChatGPT with visual models for multimodal undertakings, especially in the Visual Question Answering (VQA) framework. BLIP-2 [27], notable for its Q-former model, has shown encouraging outcomes in VQA tasks, both in zero-shot and fine-tuning settings. LLaMA-Adapter [11] enhances multimodal fine-tuning efficiency by integrating adaptation prompt vectors as adjustable parameters, showcasing versatility in multimodal contexts. MiniGPT-4 [71], derived from GPT-4 and incorporating elements from BLIP-2 and Vicuna [5], specializes in caption generation and model refinement through image-text pair fine-tuning. LLaVA [31], leveraging GPT-4, focuses on a broad spectrum of instruction fine-tuning data, ranging from multi-turn QA to image descriptions, adopting a dual-stage fine-tuning approach that prioritizes language model loss while keeping the visual model static.
2.2 Federated Learning with Heterogeneous Data
@@
Current methodologies tackling the challenge of data heterogeneity fall into the following broad categories. Some approaches aim to simultaneously improve the models on both clients and server sides through optimization techniques. Key contributions in this area have been made by the work of [26, 17, 18, 10], who have investigated various optimization methods. Other strategies focus on enhancing the stability of local models via knowledge transfer, a technique that is model-agnostic and has been explored in the research by [55, 61, 4], aiming to mitigate data heterogeneity by spreading local training knowledge throughout the whole FL framework. Additional methods, such as those proposed by [30, 2], concentrate on improving model aggregation on the server side to address data heterogeneity. Certain strategies also regulate the scheduling of client participation to avoid biasing the FL model towards classes that are more prevalent, as explored in the studies by [56, 63]. While these approaches have advanced the handling of data heterogeneity, they often do not fully address the specific issues related to long-tailed distributions in FL.
@@
A recent approach, CReFF [44], introduces a decoupling strategy to create balanced class-distribution federated features for the server model and to retrain the classifier with these features. Nonetheless, CReFF encounters two main limitations due to its reliance on generating federated features through client-side gradient information: 1) The one-to-many relationship between gradients and samples can result in the problem becoming ill-posed; 2) The absence of semantic guidance might lead to federated features that lack discriminative ability for their respective classes. The subsequent attempt, CLIP2FL [45], seeks to overcome these drawbacks by integrating a multimodal model to direct the federated learning process. However, it still has its own drawbacks. Firstly, deploying the sizable CLIP model on devices increases memory and computational demands. Secondly, transmitting both the gradient and parameters of local models to the server, as necessitated by both CLIP4FL and CReFF, raises significant privacy concerns, as attackers could potentially reconstruct client images through reverse engineering [12, 72, 13].
3 Methodology
@@
In the conventional federated learning (FL) pipeline, the FL models are typically assigned to local clients for training on their heterogeneous datasets. These models are then sent back to the server for aggregation. This cycle continues repeatedly until the FL training concludes. To better align the above FL framework with practical requirements, we incorporate the following two additional stages. The first stage occurs before local training, involving pretraining on the server side, a strategy supported by previous work [3, 49] which found that pretraining can accelerate the convergence of FL training and mitigate the effects of data heterogeneity on convergence. The second stage takes place after aggregation, where the FL model may undergo further training to meet the broader requirements of FL companies, such as the performance and safety considerations we discuss later.
@@
Drawing inspiration from recent works in learning paradigms of NLP [40, 8, 62, 52, 65, 25, 21, 49, 59], we structure our work into three parts: global multimodal pretraining, federated local finetuning, and global alignment. We deploy multimodal approaches at the server side to assist both global multimodal pretraining and global alignment phases. In this section, we will introduce our comprehensive framework as follows: Section 3.1 will delve into the details of our global multimodal pretraining; Section 3.2 will cover federated local finetuning; Section 3.3 will discuss global alignment; and in Section 3.4, we will compare our method with previous approaches.
3.1 Global Multimodal Pretraining
@@
As discussed in Sec. 1, the wealth of open-source multimodal data, such as images and their captions, remains underutilized resources for pretraining FL models. Often, these datasets are noisy, unlabeled or contain elements that are too complex, making them unsuitable for straightforward pretraining of the compact FL models. However, given the current advanced capabilities in multimodal processing of MLLMs, we now have new, convenient methods to leverage such data for pretraining purposes. Utilizing GPT-4, akin to the approach used in LLaVA, we can transform complex image data collected from the internet into three main categories:
@@
Figure 2: The visualization of our pretraining mechanism
Our pretraining mechanism draws inspiration from the structure of LLaVA, a highly effective and recent multimodal large language model. It consists of three key components: a visual encoder $g$, which is a frozen pretrained CLIP model; a projection layer designed to align the features of the visual model with the text domain embeddings, where the projection layer is a trainable matrix $W$; and a part comprising a large language model (LLM), typically employing promising models such as Vicuna or LLaMA-2. The workflow of LLaVA proceeds as follows: For the data formats mentioned earlier, whether it be a conversation, detailed description, or complex reasoning, the input includes a text modality instruction $\mathbf{X}_{\mathrm{q}}$ (e.g., "Could you provide a detailed description of this image?") and an image $\mathbf{X}_{\mathrm{v}}$. The instruction $\mathbf{X}_{\mathrm{q}}$ passes through an embedding layer to obtain the text embedding $\mathbf{H}_{q}$. As for the image $\mathbf{X}_{\mathrm{v}}$, it first goes through the visual encoder $g$ to acquire the grid feature $\mathbf{Z}_{\mathrm{v}}$, which then passes through the projection layer to obtain the visual embedding $\mathbf{H}_{v}$, aligned in dimension with $\mathbf{H}_{q}$:
$\mathbf{H}_{v}=\mathbf{W}\cdot\mathbf{Z}_{\mathrm{v}}\text{, where }\mathbf{Z}_{\mathrm{v}}=g\left(\mathbf{X}_{\mathrm{v}}\right)$ (1)
@@
Drawing from the LLaVA mechanism, our Global Multimodal Pretraining essentially integrates the compact FL model, designed for downstream image classification tasks, as a component within the LLaVA framework’s visual encoder. The FL model is made trainable and is denoted as $g_{fl}$. Inspired by the previous work in knowledge distillation[15, 38, 20, 64] and Bayesian Algorightms [68, 1, 53, 69, 67, 32, 70], We have developed an approach termed Dynamic Weighted Distillation, which involves computing a weighted average of the visual features obtained from the FL model and those from the original visual encoder:
$\mathbf{Z}_{\mathrm{v}}=(1-\alpha)g\left(\mathbf{X}_{\mathrm{v}}\right)+\alpha g_{fl}\left(\mathbf{X}_{\mathrm{v}}\right)$ (2)